OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12

OpenStudy (anonymous):

2x − y = 6 4x − 2y = 12 -2(2x − y = 6) 4x − 2y = 12 -4x + 2y = -12 4x − 2y = 12 0 = 0 THis is an identity situation. Identity situations mean that it's the same line and are therefore infinitely many solutions.

OpenStudy (anonymous):

That means it can be any number right?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

X can be any number I mean

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):